OpenStudy (anonymous):

What is the simplified form of the expression (4x^3)^3 ? pronounced like four X. Raised to the 3rd power, raised to the 3rd power.

jimthompson5910 (jim_thompson5910):

you use the rule that (a*b)^c = a^c * b^c and the rule (a^b)^c = a^(b*c) to get (4x^3)^3 = (4)^3 * (x^3)^3 (4x^3)^3 = 64 * (x^3)^3 (4x^3)^3 = 64 * x^(3*3) (4x^3)^3 = 64x^9

OpenStudy (anonymous):

Thanks Jim!

jimthompson5910 (jim_thompson5910):

you're welcome